# Preference-Based Gradient Estimation for ML-Based Approximate Combinatorial Optimization

### v1

@@ Abstract
CO problems arise in a wide range of fields from medicine to logistics and manufacturing. While exact solutions are often not necessary, many applications require finding high-quality solutions quickly. For this purpose, we propose a data-driven approach to improve existing non-learned approximation algorithms for combinatorial optimization (CO). We parameterize the approximation algorithm and train a graph neural network (GNN) to predict parameter values that lead to the best possible solutions. Our pipeline is trained end-to-end in a self-supervised fashion using gradient estimation, treating the approximation algorithm as a black box. We propose a novel gradient estimation scheme for this purpose, which we call preference-based gradient estimation. Our approach combines the benefits of the neural network and the non-learned approximation algorithm: The GNN leverages the information from the dataset to allow the approximation algorithm to find better solutions, while the approximation algorithm guarantees that the solution is feasible. We validate our approach on two well-known combinatorial optimization problems, the travelling salesman problem and the minimum $k$-cut problem, and show that our method is competitive with state of the art learned CO solvers.
Combinatorial Optimization, Combinatorial Optimisation, Gradient Estimation, Preference Learning, Graph Neural Networks, Travelling Salesman Problem, Traveling Salesman Problem, TSP, Minimum k-Cut, Machine Learning, ICML
@@ 1 Introduction
In order to avoid such drawbacks, we instead use the GNN to augment an existing, non-learned approximation algorithm for a given CO problem. We parameterize the approximation algorithm and let the GNN estimate the parameters based on the input graph. By selecting an approximation algorithm that is fast enough to be used during training, we can use the same overall pipeline during training and testing, addressing the concern raised by Xia et al. (2024). Our approach combines the advantages of the GNN and the traditional, non-learned approximation algorithm, and allows them to cancel out each other’s weaknesses. The GNN leverages a data-driven approach to achieve good performance on the problem instances that occur frequently. However, it cannot output solutions to the CO problem directly. The CO approximation algorithm solves this by transforming the parameters predicted by the GNN into a feasible solution. On its own, the approximation algorithm cannot leverage the information in the dataset and it therefore performs comparatively poorly in practice, but this is offset by the GNN.
Since the CO approximation algorithm’s output is discrete, the algorithm is not differentiable. We therefore use gradient estimation to allow us to backpropagate through the approximation algorithm during training. We experiment using several existing gradient estimators, but we also introduce a new gradient estimation scheme, which improves over existing approaches for this purpose. We call our new scheme preference-based gradient estimation (PBGE). Our training setup allows for fully self-supervised training without the use of pre-computed optimal solutions to the CO problems.
@@
  3. 3.
    Extensive experimental evaluation of (1) and (2) on two common CO problems.
Our implementation and data will be made public as soon as the paper is accepted [^1] The code was submitted as supplementary material for the reviewers and will be published once the paper is accepted.
@@ 2 Related Work
The easiest way of training a model for CO is to assume the existence of ground truth solutions to the CO problems and train in a supervised fashion. Vinyals et al. (2015) leverage the fact that many CO problems ask to identify a certain subset of the input (e.g. minimum-k-cut) or a permutation of the input (e.g.  travelling salesman problem (TSP)). They introduce an LSTM-based sequence-to-sequence architecture where the elements of the output sequence are positions in the input sequence, which they name pointer network. Nowak et al. (2017); Joshi et al. (2019) recognize that most common CO problems have a natural graph representation. In their work, a GNN predicts an approximate solution as a heatmap, which is then decoded into a valid solution to the CO problem using beam search. Georgiev et al. (2023) follow a neural algorithmic reasoning approach to learn to imitate CO solvers. The model is pre-trained on simple algorithms and then fine-tuned on difficult ones. Finally, Sun & Yang (2023) uses graph-based denoising diffusion to generate high-quality solutions. However, these supervised approaches aren’t applicable to such cases where calculating exact solutions for the training problems is not feasible.
@@
Several approaches have used reinforcement learning (RL) to remove this dependence on a labeled dataset. Bello et al. (2017) use a pointer network (Vinyals et al., 2015), but train it with RL. Khalil et al. (2017); Kool et al. (2019) use the GNN autoregressively to predict which node should be added next to the solution set and repeat that process until a valid solution is reached. Xu et al. (2020) formulate the CO problem as a Markov decision process, then use an algorithm similar to AlphaGo Zero (Silver et al., 2017) to solve it autoregressively. Kim et al. (2024) propose a modification to the RL training process for autoregressive methods that improves sample efficiency. However, the use of the score function estimator in many of these methods leads to high-variance gradients, which makes training more difficult.
Other self-supervised approaches that do not rely on reinforcement learning include Duan et al. (2022), which use a contrastive loss instead. Schuetz et al. (2022) focus on quadratic and polynomial unconstrained binary optimization, which many CO problems can be formulated as. This allows them to make use of a self-supervised loss function specific to these two problem families. Sanokowski et al. (2024) also concentrate on quadratic unconstrained binary optimization, but employ a diffusion-based approach similar to Sun & Yang (2023). Similarly, Karalias & Loukas (2020); Bu et al. (2024) formulate a self-supervised loss function for a comprehensive class of CO problems. Toenshoff et al. (2021) use an LSTM-based architecture to solve binary maximum constraint satisfaction problems, which many CO problems can be formulated as. Since these approaches focus on certain families of CO problems, they aren’t general to CO as a whole.
@@
Gasse et al. (2019) use a GNN as a heuristic for the branch-and-bound algorithm, which guarantees exact solutions at the expense of a comparatively longer running time. The GNN is trained via imitation learning from a known high-quality but slow heuristic. Joshi et al. (2021) compare some of the paradigms introduced in other papers in structured experiments. There have been two lines of work to address the problem of backpropagating through combinatorial optimization problems. Firstly, if we have a set of optimal solutions given as training data, we can use supervised learning to train the GNN to output adjacency matrices as close as possible to the optimal solutions (Elmachtoub & Grigas, 2022). This is often called “predict, then optimize”. Secondly, there are several methods to backpropagate through a non-differentiable CO algorithm, such as Niepert et al. (2021); Minervini et al. (2023); Vlastelica et al. (2020). Related to our work is decision-focused learning, which has developed several methods to backpropagate through CO solvers (Mandi et al., 2023).
@@ 3 Background
3.1 Combinatorial Optimization Problems
A combinatorial optimization (CO) problem asks us, given a discrete set $M$ and an objective function $J_{\text{CO}}:M\longrightarrow\mathbb{R}$, to find the minimum
@@ 3.1 Combinatorial Optimization Problems
Of course, maximization problems can be turned into minimization problems by inverting the sign. Since finding the exact global optimum is often not necessary in practice, this paper focuses on finding approximate solutions.
To illustrate our approach, we will refer to specific problems in combinatorial optimization as follows.
Minimum $k$-cut problem.
We are given a connected, undirected graph ${\mathcal{G}}=(V,E,w)$ with edge weights $w:E\longrightarrow\mathbb{R}_{>0}$, as well as a desired number of clusters $k\in\mathbb{N},2\leq k\leq|V|$. The goal is to find a set of edges $C\subseteq E$ with minimal total weight whose removal leaves $k$ connected components. This set is called a minimum $k$-cut.
Formally, we are optimizing
$\min_{C\subseteq E}\sum_{e\in C}w(e),$
[…]
A commonly used approximation algorithm for solving the minimum $k$-cut problem is the Karger–Stein algorithm (Karger & Stein, 1993), which we describe in appendix A.1.
[…]
The version of the TSP most commonly experimented on in related literature (Kool et al., 2019; Joshi et al., 2019, 2021) is called Euclidean TSP. We are given a complete, undirected graph ${\mathcal{G}}=(V,E,w)$ whose nodes $V$ lie at points in the unit square, and whose edge weights $w$ are the Euclidean distances between the respective points. The goal is to find a minimum-weight Hamiltonian cycle, i.e. a cycle that visits every node exactly once and where the sum of the weights of edges that are traversed in the cycle is minimal.
A well-known probabilistic approximation algorithm for solving the TSP is the random insertion algorithm (Karg & Thompson, 1964), which we describe in appendix A.2.
3.2 Residual Gated Graph Convnets
We use residual gated graph convnets (Bresson & Laurent, 2017), but adapt them to include edge features $\bm{e}_{ij}^{l}$ and a dense attention map $\bm{\eta}_{ij}^{l}$ following (Joshi et al., 2019). The input node features $\bm{x}_{i}^{0}$ and edge features $\bm{e}_{ij}^{0}$ are first pre-processed using a single-layer MLP for each of the two. Each further layer is computed as follows:
$\bm{x}_{i}^{l+1}$ $=\bm{x}_{i}^{l}+\mathrm{ReLU}\Bigg{(}\mathrm{BN}\bigg{(}W_{1}^{l}\bm{x}_{i}^{l}+\sum_{j\in\mathcal{N}_{i}}\bm{\eta}_{ij}^{l}\odot W_{2}^{l}\bm{x}_{j}^{l}\bigg{)}\Bigg{)}$
[…]
The final edge-level output is calculated from the last layer’s edge features $\bm{e}_{ij}^{l}$ using another MLP. For the remainder of the paper, $f({\mathcal{G}})\in\mathbb{R}^{|E|}$ refers to applying this GNN on a graph ${\mathcal{G}}$.
3.3 REINFORCE
Given a function $J$ and a parameterized probability distribution $p_{\theta}(x)$, the REINFORCE algorithm (Williams, 1992), also known as the score function estimator, estimates the true gradient as follows:
$\approx\frac{1}{S}\sum_{i=1}^{S}J(y_{i})\nabla_{\theta}\log p_{\theta}(x_{i}),\quad y_{i}\sim p_{\theta}(x_{i})$
@@ 3.3 REINFORCE
While this estimator is applicable even for discrete distributions and is unbiased, it suffers from large variance. implicit maximum likelihood estimator (I-MLE) (Niepert et al., 2021) is another gradient estimator that produces biased estimates with smaller variance.
Figure 1: Overview of the training process. In the forward pass, the GNN produces edge scores that are used as parameters for a CO approximation algorithm. The approximation algorithm operates on these parameters by using them to scale the graph’s edge weights and running an off-the-shelf approximation algorithm on the modified graph to obtain a solution $\hat{\bm{y}}$. Since the CO approximation algorithm is not differentiable in general, a gradient estimation scheme is used in the backward pass. We propose a novel scheme for this purpose, preference-based gradient estimation (PBGE). Note that there is no explicit loss function, since the gradient is estimated directly from the solution $\hat{\bm{y}}$.
@@ 4 Problem Statement
We consider CO problems on graphs with a linear objective function $J_{\text{CO}}$ and a probabilistic approximation algorithm $h(\hat{\bm{y}}\mid{\mathcal{G}})$. The approximation algorithm takes as input a graph ${\mathcal{G}}=(V,E,w)$ with nodes $V$, edges $E$ and edge weights $w:E\longrightarrow\mathbb{R}_{>0}$, and returns (samples) a potentially suboptimal solution $\hat{\bm{y}}$. For instance, we might have the minimum $k$-cut problem as defined in subsection 3.1.
We now want to use a GNN $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ applied to the input graphs ${\mathcal{G}}$ to compute an updated graph ${\mathcal{G}}^{\prime}=f_{\bm{\theta}}({\mathcal{G}})$ such that the probabilistic approximation algorithm when applied to this new graph is improved in expectation.
Hence, we want to solve the following optimization problem:
@@
For each input graph ${\mathcal{G}}$, $h(\hat{\bm{y}}\mid f_{\bm{\theta}}({\mathcal{G}}))$ is a discrete probability distribution (due to the assumption that $h$ is probabilistic) parameterized by $\bm{\theta}$. The main challenge for solving this optimization problem is that the (discrete) approximation algorithms are typically not differentiable functions and that optimal solutions are prohibitively expensive to obtain as training data. Moreover, we assume that the approximation algorithm is a black box. This means that while we can sample from the probability distribution defined by it, we cannot calculate a probability for a given sample.
Note that if the approximation algorithm available to us is deterministic, we can convert it into a probabilistic approximation algorithm by simply adding noise to the input, as demonstrated by e.g. Niepert et al. (2021). For instance, this can take the form of adding noise from a Gumbel distribution to the edge weights of the input graph.
With this paper, we propose PBGE, a new approach based on preference-based learning. It estimates gradients by contrasting better and worse solutions, similar to preference-based learning in the context of large language models.

### v2

# Preference-Based Gradient Estimation for ML-Guided Approximate Combinatorial Optimization

@@ Abstract
CO problems arise across a broad spectrum of domains, including medicine, logistics, and manufacturing. While exact solutions are often computationally infeasible, many practical applications require high-quality solutions within a given time budget. To address this, we propose a learning-based approach that enhances existing non-learned approximation algorithms for combinatorial optimization (CO). Specifically, we parameterize these approximation algorithms and train graph neural networks to predict parameter values that yield near-optimal solutions. Our method is trained end-to-end in a self-supervised fashion, using a novel gradient estimation scheme that treats the approximation algorithm as a black box. This approach combines the strengths of learning and traditional algorithms: the GNN learns from data to guide the algorithm toward better solutions, while the approximation algorithm ensures feasibility. We validate our method on two well-known combinatorial optimization problems: the travelling salesman problem (TSP) and the minimum $k$-cut problem. Our results demonstrate that the proposed approach is competitive with state-of-the-art learned CO solvers.
@@ 1 Introduction
To overcome these limitations, we propose to use GNNs to augment highly efficient approximation algorithms that are commonly available for many CO problems. The GNNs are trained to predict parameters for the approximation algorithms that influence their behavior. Specifically, we predict edge weights for the input graphs and then apply the approximation algorithms to this modified graph. By choosing fast approximation algorithms, we can incorporate them into the training loop and use the exact same pipeline during training and inference. Our method combines the strengths of data-driven and established heuristic algorithms: the GNNs learn to perform well on frequently encountered instances, and the approximation algorithm, in turn, ensures the feasibility of the resulting solutions.
Once trained, the GNN-augmented approximation algorithms serve as drop-in replacements, preserving efficiency and usability while improving solution quality. Indeed, we provide a theoretical guarantee that, for a broad class of approximation algorithms, a suitably trained upstream GNN can learn parameters that render the approximation algorithm optimal.
Since the approximation algorithms return discrete solutions, the gradients of their outputs with respect to their inputs are zero almost everywhere. We therefore need to apply gradient estimation techniques such as the score function and straight-through estimators (Williams, 1992; Bengio et al., 2013) to backpropagate through the algorithms. Since existing gradient estimators cannot be used in a self-supervised setting or have high variance, we propose preference-based gradient estimation (PBGE), which estimates gradients through a comparison of solutions sampled from the given approximation algorithm. This approach enables fully self-supervised training, eliminating the need to pre-compute exact solutions as ground-truth labels, an often costly or even infeasible requirement for many CO problems. Figure 1 illustrates the proposed PBGE framework.
Figure 1: Overview of the proposed framework for ML-guided approximate combinatorial optimization. In the forward pass, the GNN predicts edge scores that are used as input parameters for a CO approximation algorithm. The approximation algorithm operates on these parameters by using them to scale the graph’s edge weights and running an off-the-shelf approximation algorithm on the modified graph to obtain a solution $\hat{\bm{y}}$. Since the CO approximation algorithm is not differentiable in general, a gradient estimation scheme is used in the backward pass.
We evaluate our approach on two well-known combinatorial optimization problems, the travelling salesman problem (TSP) and the minimum $k$-cut problem. Our approach improves the performance of a commonly used approximation algorithm by an order of magnitude in the case of minimum $k$-cut while only minimally increasing the runtime.
@@
    A method for augmenting existing approximation algorithms for CO using GNNs;
  2. 2.
@@
  4. 4.
    An extensive experimental evaluation of (1), (2), and (3) on common CO problems.
The code is part of the supplementary material and will be made public upon acceptance of the paper.
@@ 2 Related work
The easiest way of training a model for CO is to assume the existence of ground truth solutions to the CO problems and train in a supervised fashion. Nowak et al. (2017); Joshi et al. (2019) use a GNN to predict an approximate solution as a heatmap, which is then decoded into a feasible solution to the CO problem using beam search. Vinyals et al. (2015) introduce pointer networks, which leverage the fact that many CO problems ask to identify a subset or a permutation of the input. Georgiev et al. (2023) follow a neural algorithmic reasoning approach to learn to imitate CO solvers. Gasse et al. (2019); Kahng et al. (2024) replace components of an existing algorithm that would otherwise be expensive to compute with neural networks. Antoniadis et al. (2025) augment existing algorithms using discrete predictions. Finally, Sun & Yang (2023); Yu et al. (2024) use graph-based denoising diffusion to generate high-quality solutions. However, these supervised approaches aren’t applicable to such cases where calculating exact solutions for the training problems is not feasible.
@@
Several approaches have used reinforcement learning (RL) to remove this dependence on a labeled dataset (Kwon et al., 2021; Qiu et al., 2022; Kim et al., 2024). A common approach is to formulate the CO problem as a Markov decision process (Bello et al., 2017; Deudon et al., 2018; Xu et al., 2020; Wu et al., 2022). Khalil et al. (2017); Kool et al. (2019) use the GNN autoregressively to predict which node should be added next to the solution set and repeat that process until a valid solution is reached. Cheng et al. (2023); Pan et al. (2023); McCarty et al. (2021) tackle large-scale instances by locally optimizing sub-parts of the instance individually. However, RL is often sample inefficient and difficult to train due to the high variance of the gradient estimations.
A common self-supervised approach that does not rely on RL is to formulate a surrogate loss consisting of a term that encourages high quality solutions and one that softly enforces the constraints (Karalias & Loukas, 2020; Sun et al., 2022; Min et al., 2022, 2023; Bu et al., 2024; Wenkel et al., 2024). Duan et al. (2022) train self-supervised using a contrastive loss instead. Schuetz et al. (2022); Sanokowski et al. (2024) focus on quadratic and polynomial unconstrained binary optimization, allowing them to formulate self-supervised loss functions for these problem families. Tönshoff et al. (2021) use an LSTM-based architecture to solve binary maximum constraint satisfaction problems, which many CO problems can be formulated as. Corsini et al. (2024) train a pointer network-based model in a self-supervised fashion. Zhang et al. (2023) tackle CO using GFlowNets (Bengio et al., 2021). Yau et al. (2024) propose a GNN architecture that can capture message passing algorithms with provable approximation guarantees for a large class of CO problems.
@@
Joshi et al. (2021) compare some of the paradigms introduced in other papers in structured experiments. Xia et al. (2024) argue that the common pattern of training the GNN using a surrogate loss, but testing it using a decoder, means that the GNN is applied inconsistently, leading to uncertain performance during testing. Our approach addresses this problem by incorporating the decoder during training. There have been two lines of work on backpropagating through CO problems. Firstly, if we have a set of optimal solutions given as training data, we can use supervised learning to train the GNN to output adjacency matrices as close as possible to the optimal solutions (Elmachtoub & Grigas, 2022). This is often called “predict, then optimize”. Secondly, there are several methods to backpropagate through a non-differentiable CO algorithm, such as Niepert et al. (2021); Minervini et al. (2023); Vlastelica et al. (2020). Related to our work is decision-focused learning, which has developed several methods to backpropagate through CO solvers (Mandi et al., 2024). Our approach follows this paradigm. The implicit maximum likelihood estimator (I-MLE) (Niepert et al., 2021) is another such method; we compare against it in our experiments and show that our approach improves over I-MLE.
@@ 3 Background
Combinatorial optimization problems.
A combinatorial optimization (CO) problem asks us, given a discrete set $M$ and an objective function $J_{\text{CO}}:M\rightarrow\mathbb{R}$, to find the minimum
@@
Maximization problems can be turned into minimization problems by inverting the sign of the objective function. Since finding the exact global optimum is often not required in practice, this paper focuses on efficiently finding approximate solutions.
To illustrate our approach, we will refer to specific CO problems as follows. The minimum $k$-cut problem asks, given a weighted, undirected graph and a desired number $k$ of connected components, to find a set of edges with minimum total weight whose removal leaves the graph with exactly $k$ connected components. A commonly used approximation algorithm for this problem is the Karger–Stein algorithm (Karger & Stein, 1993). In the travelling salesman problem (TSP), we are given a weighted, undirected graph, and are asked to find a minimum-weight Hamiltonian cycle, i.e. a cycle that visits every node exactly once and where the sum of the weights of edges that are traversed in the cycle is minimal. A well-known probabilistic approximation algorithm for solving the TSP is the random insertion algorithm (Karg & Thompson, 1964). The variant of the TSP most commonly experimented on in related literature (Kool et al., 2019; Joshi et al., 2019, 2021) is the Euclidean TSP, where the graph is fully connected, the nodes represent points in the unit square and the edge weights are the distances between these points. We give formal definitions for both problems as well as descriptions of the algorithms mentioned in Appendix A.
Graph neural networks (GNNs).
Let $\mathcal{G}=(V,E)$ be a graph with $n=|V|$ the number of nodes. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the feature matrix that associates each node of the graph with a $d$-dimensional feature vector and let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be the adjacency matrix. GNNs based on the message passing paradigm have three basic operations, which are defined as
$\mathbf{h}_{i}^{\ell}=\gamma\left(\mathbf{h}_{i}^{\ell-1},\varphi_{j\in\mathcal{N}(v_{i})}\phi\left(\mathbf{h}_{i}^{\ell-1},\mathbf{h}_{j}^{\ell-1},r_{ij}\right)\right),$
where $\gamma$, $\varphi$, and $\phi$ represent update, aggregation and message function respectively.
Propagation step. The message-passing network computes a message $m_{ij}^{\ell}=\phi(\mathbf{h}_{i}^{\ell-1},\mathbf{h}_{j}^{\ell-1},r_{ij})$ between every pair of neighboring nodes $(v_{i},v_{j})$. The function takes in input $v_{i}$’s and $v_{j}$’s representations $\mathbf{h}_{i}^{\ell-1}$ and $\mathbf{h}_{j}^{\ell-1}$ at the previous layer $\ell-1$, and the relation $r_{ij}$ between the two nodes.
Aggregation step. For each node in the graph, the network performs an aggregation computation over the messages from $v_{i}$’s neighborhood $\mathcal{N}(v_{i})$ to calculate an aggregated message $M_{i}^{\ell}=\varphi(\{m_{ij}^{\ell}\mid v_{j}\in\mathcal{N}(v_{i})\})$. The definition of the aggregation function differs between methods.
Update step. Finally, the model non-linearly transforms the aggregated message $M_{i}^{\ell}$ and $v_{i}$’s representation from previous layer $\mathbf{h}_{i}^{\ell-1}$ to obtain $v_{i}$’s representation at layer $\ell$ as $\mathbf{h}_{i}^{\ell}=\gamma(\mathbf{h}_{i}^{\ell-1},M_{i}^{\ell})$.
Gradient Estimators for Discrete Distributions.
Learning systems that backpropagate through discrete distributions and algorithms require gradient estimation approaches. This is because gradients cannot be directly computed or are zero almost everywhere for operations such as sampling from a categorical distribution or running a solver for CO problems. The standard estimator is the score function estimator, also known as the REINFORCE estimator (Williams, 1992). Given a function $J$ and a parameterized probability distribution $p_{\theta}(x)$ it estimates the true gradient as follows:
$\nabla_{\theta}\mathbb{E}_{y\sim p_{\theta}(x)}\big{[}J(y)\big{]}=\mathbb{E}_{y\sim p_{\theta}(x)}\big{[}J(y)\nabla_{\theta}\log p_{\theta}(x)\big{]}\approx\frac{1}{S}\sum_{i=1}^{S}J(y_{i})\nabla_{\theta}\log p_{\theta}(x_{i}),\quad y_{i}\sim p_{\theta}(x_{i})$
@@
While this estimator is unbiased, it suffers from large variance. There are other estimators, such as the straight-through estimator (Bengio et al., 2013) and I-MLE (Niepert et al., 2021) with smaller variance but biased gradient estimates.
@@ 4 Problem statement
We consider CO problems on graphs with a linear objective function $J_{\text{CO}}$ and a probabilistic approximation algorithm $h(\hat{\bm{y}}\mid{\mathcal{G}})$. The approximation algorithm takes as input a graph ${\mathcal{G}}=(V,E,w)$ with nodes $V$, edges $E$ and edge weights $w:E\rightarrow\mathbb{R}_{>0}$, and returns (samples) a potentially suboptimal solution $\hat{\bm{y}}$. For instance, we might have the minimum $k$-cut problem and the Karger–Stein algorithm.
We now want to use a GNN $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ applied to the input graphs ${\mathcal{G}}$ to compute an updated graph ${\mathcal{G}}^{\prime}=f_{\bm{\theta}}({\mathcal{G}})$ such that the probabilistic approximation algorithm when applied to this new graph is improved in expectation. Hence, we want to solve the following optimization problem:
@@
For each input graph ${\mathcal{G}}$, $h(\hat{\bm{y}}\mid f_{\bm{\theta}}({\mathcal{G}}))$ is a discrete probability distribution (due to the assumption that $h$ is probabilistic) parameterized by $\bm{\theta}$. The main challenge in solving this optimization problem is that (discrete) approximation algorithms are typically not differentiable functions and that optimal solutions are prohibitively expensive to obtain as training data, making supervised training infeasible. Moreover, we assume that the approximation algorithm is a black box—while we can sample from the probability distribution defined by it, we cannot compute a probability mass for a given sample.